OpenStudy (anonymous):

Let ρAl represent the density of aluminum and ρFe that of iron. Find the radius of a solid aluminum sphere that balances a solid iron sphere of radius rFe on an equal-arm balance. (Use any variable or symbol stated above as necessary.)

OpenStudy (anonymous):

does anyone know how to get this?

OpenStudy (aravindg):

:)

OpenStudy (saifoo.khan):

it's a bit tricky.

OpenStudy (saifoo.khan):

deals with the formula: Density = Mass/Volume

OpenStudy (anonymous):

im so confused with it..decided to come here and get some help lol

OpenStudy (saifoo.khan):

LOL. sorry i have no idea.

OpenStudy (anonymous):

i'm just playing with this in my head at the moment, and i could be completely wrong but i'm thinking; since density = mass/volume the weight for both spheres has to be the same (to balance) so the mass will be equal. the density is just a ratio of mass to volume so if you divided pFe by pAl you might be able to find the ratio needed in volume. so my working came out to be something like... \[rAl = (pFe/pAl) * rFe\] someone care to explain/correct?

OpenStudy (anonymous):

and then obviously reverse the sphere volume equation to find radius from a given volume

OpenStudy (anonymous):

im not sure..thats all that is given to me to figure out

OpenStudy (anonymous):

ok i thought it over a bit more and i made a mistake with my previous calc i think. basically for the weight to balance both sides need to be equal weight, so massFe will equal massAl. so in the \[pFe = mFe / vFe\]equation mFe = mAl so you can sub in\[pFe = (pAl * vAl) / vFe\] then rearrange to solve for vAl and then use the volume of aluminium in the sphere volume equation \[vAl = 4/3 * r^3\] to solve for the radius of the aluminium sphere. It's a bit tricky with no numbers (I had to put my own numbers in to check) but I THINK this works. Maybe an actual physics student can correct me.

OpenStudy (anonymous):

so you think the answer is 4/3* r^3?

OpenStudy (anonymous):

There may be an easier way to do it because the way I suggested is a little convoluted, but if I had to put an answer down I'd put \[vAl = pAl*vFe/pAl\] and then \[3(vAl)/4*\pi=rAl^3\]

OpenStudy (anonymous):

unfortunately that isnt the answer..and the answer has still not been given to me